### v5

Abstract
The rapid growth of preprint servers has accelerated scientific dissemination but has also shifted the technical burden of manuscript preparation to authors. This challenge is particularly acute in computational research, where manuscripts must remain synchronised with evolving data and code. We present Rxiv-Maker, a framework that resolves this by converting simple Markdown files into professionally typeset, publication-ready PDFs. Its core feature is the ability to execute embedded code, creating a self-updating manuscript where figures and statistical values are generated directly from source data during compilation. This ensures that the final document is always current and fully reproducible. By integrating with standard tools like Git and Visual Studio (VS) Code, Rxiv-Maker provides an efficient, transparent, and collaborative authoring experience, applying principles of software engineering to academic writing to foster open and verifiable science.
keywords:
article template | scientific publishing | preprints
{corrauthor}
(B. M. Saraiva) bsaraiva and itqb.unl.pt; (G. Jaquemet) guillaume.jacquemet and abo.fi; (R. Henriques) ricardo.henriques and itqb.unl.pt
Introduction
The landscape of scientific publishing has been profoundly reshaped by the rise of preprint servers, which facilitate rapid dissemination of research findings across numerous platforms beck2020; levchenk2024; Fraser2020_preprint_growth (Fig. Methods, Fig. Methods). This acceleration, however, has transferred the complex task of typesetting from publishers to researchers Vale2015_preprints; Tenant2016_academic_publishing; lin2020. For those of us in computational fields, this is compounded by a more pressing challenge: ensuring the manuscript remains perfectly synchronised with our data and analysis code. We have all faced the tedious and error-prone task of manually updating a p-value or a sample size in the text after re-running an analysis perkel2022. This disconnect between the research and the report is a critical issue, particularly in disciplines like bioimage analysis, where our findings are built upon complex computational pipelines biaflows2024; dl4miceverywhere2024.
Figure 1: System Architecture. Rxiv-Maker integrates Markdown content, YAML metadata, executable scripts, and bibliographies through a processing engine that combines local execution with LaTeX PDF compilation or Word (.docx) generation to produce publication-ready documents.
Figure 2: Processing Pipeline. User-provided content (left) undergoes automated processing (right), including parsing, script execution, document compilation (PDF/Word), and final output generation.
To address these challenges, we developed Rxiv-Maker, a framework designed to streamline the creation of scientific manuscripts. It automates the production of publication-ready documents from simple Markdown, harnessing the typesetting power of LaTeX or the flexibility of Microsoft Word without requiring the author to manage complex formatting manually. Rxiv-Maker transforms the document into a self-updating manuscript by executing embedded Python or R scripts during compilation. This approach forges a direct, traceable link between the data and the final publication, eliminating manual transcription errors and ensuring the document is always synchronised with the latest results. The framework is designed as a local-first system, providing a rapid and responsive authoring experience that integrates seamlessly with version control systems like Git, thereby enhancing transparency and collaboration Ram2013_git_science; Perez-Riverol2016_github_bioinformatics. The overall architecture and processing pipeline are illustrated in Fig. 1 and Fig. 2, respectively. A tutorial is provided to guide new users through their first manuscript (Methods).
Results
0.0.1 From Static Text to Self-Updating Manuscripts
A central achievement of Rxiv-Maker is its ability to create dynamic, self-updating documents. By embedding executable code snippets directly within the Markdown source, the manuscript transitions from a static report to a living document. This functionality enables the direct execution of analysis scripts during compilation, with the results (whether statistical values, tables, or figures) directly injected into the text. This process eliminates the possibility of transcription errors when, for example, copying a value from a terminal or another program into the manuscript. We demonstrate this with a practical example from this very paper. The Python code shown in Fig. 3 is embedded in the source of our manuscript. At build time, it retrieves the latest public data on arXiv submissions, calculates key statistics, and inserts them into the text. This ensures that our discussion of preprint growth is always based on the most current evidence available. The analysis reveals that, since 1991, a total of 2.91 million submissions have been made over 35 years (Fig. Methods). These are not static numbers; they were computed when this document was compiled on December 17, 2025, using data updated in December 2025. This capability extends to all forms of programmatic figure generation, as detailed in Methods, ensuring that every visual and numerical claim is reproducibly generated.
Figure 3: Embedded Python for dynamic content. This figure shows a Python script embedded in the manuscript’s source code. The script is executed at build time to compute data attributes (e.g., total submissions and span of years) and inject the resulting values directly into the text. This process, rendered here with syntax highlighting from the Rxiv-Maker VS Code extension, eliminates manual transcription errors and ensures the text remains synchronised with the source data.
0.0.2 Multi-Format Production: From PDF to Word
Rxiv-Maker allows researchers to achieve the typographic quality of LaTeX while writing in simple, intuitive Markdown. This is accomplished through a sophisticated translator that converts Markdown to restricted LaTeX in a guarded manner. We recognised that a naive, single-pass conversion would be too fragile for academic documents, where complex syntax for mathematics, citations, and cross-references must be preserved. Our translator first identifies and protects delicate elements: mathematical expressions (Methods), code blocks, and citation keys. Subsequent passes then normalise document structure and intelligently convert Rxiv-Maker’s extended Markdown syntax into the corresponding LaTeX commands, supporting comprehensive text formatting.
Importantly, Rxiv-Maker extends beyond PDF generation to support the collaborative realities of modern science, where co-authors or journals may require editable formats. The framework includes a dedicated Word export engine (built on python-docx) that maps the same semantic Markdown source to .docx documents. This ensures that the generated Word files maintain figure numbering, citations, and equations, allowing for seamless interchange with collaborators. Furthermore, to facilitate the review process, the track-changes feature leverages Git history to generate annotated PDFs (using latexdiff), enabling authors to visually visualize text changes between manuscript versions or against specific milestones (e.g., submission vs. revision).
The system handles complex nested formatting, inline DOI resolution, and transforms image syntax into floating figure environments, automatically handling captions, labels, and layout parameters. This enables authors to specify figure widths or positions using simple options in Markdown, without needing to modify LaTeX code. The system supports a wide array of figure generation methods, from static images to script-based visualisations (Table Methods), ensuring all visual elements are correctly and professionally typeset. For more advanced formatting needs, including complex tables, users can directly inject LaTeX code using {{tex:...}} blocks. A complete syntax reference and example are detailed in Methods.
0.0.3 An Integrated and Rapid Authoring Experience
We designed Rxiv-Maker to provide a seamless and efficient writing experience. The framework operates as a local-first command-line tool, which means that compilation is fast and authors receive immediate feedback on their changes. This rapid iteration is enhanced by an intelligent caching system that remembers which figures and results are already built. It computes a unique signature: a checksum from a figure’s source code and data dependencies. A script is only re-executed if this signature has changed, saving valuable time during compilation without compromising reproducibility, as further explained in Methods. To lower the barrier to entry, we developed a companion extension for Visual Studio Code (Fig. Methods). This extension acts as a helpful assistant, providing syntax highlighting for our extended Markdown, autocompletion for citation keys and cross-references, and real-time validation of the manuscript’s configuration. This integration of authoring and validation into a single, familiar environment streamlines the writing process, allowing researchers to focus on their content rather than on the technical details of typesetting.
0.0.4 Designed for Transparent and Collaborative Science
Reproducibility and collaboration are at the heart of Rxiv-Maker. By design, the entire manuscript is structured to be managed with Git, the version control system that is standard in software development. This includes the text, code, data, and configuration. This treats the manuscript as a complete, self-contained project where every change is tracked, attributed, and auditable. This aligns perfectly with the principles of open science, creating a transparent history of the research from its inception to the final publication. The framework’s command-line interface includes a powerful rxiv validate command that serves as a quality control gatekeeper. It performs a series of checks, such as verifying that all figures are present, all cross-references are valid, and all bibliographic entries are correctly formatted. This validation can be integrated into automated workflows, for instance, as a pre-commit hook to prevent broken versions from entering the project history, or as a check in a continuous integration (CI) pipeline to ensure that a manuscript is always ready for dissemination. This brings the rigour of professional software engineering to the academic writing process.
Discussion
Rxiv-Maker finds its place in a growing ecosystem of tools designed to enhance scientific publishing. Our approach is distinct from collaborative web-based editors like Overleaf Overleaf2024, which, while excellent for team-based LaTeX writing, do not address the fundamental challenge of keeping a manuscript synchronised with external data and analysis code HenriquesLab2015_template. Our philosophy is more closely aligned with that of platforms like Manubot himmelstein2019, Jupyter Book JupyterBook2020, and Quarto Quarto2024, which also integrate code and narrative, often building on notebook-based formats Jupyter2016_notebook. However, Rxiv-Maker is specifically optimised for the production of high-quality, submission-ready manuscripts from a local-first, developer-centric workflow. Beyond preprint distribution, rxiv-maker excels at generating professional manuscripts suitable for first submission to peer-reviewed journals, providing publication-quality PDFs and clean LaTeX source code that many journals readily accept prior to their final journal-specific formatting requirements. This positions rxiv-maker as a comprehensive solution spanning the entire academic publishing pipeline—from initial draft through preprint distribution to journal submission. As shown in our comparison with other tools (Table Methods), its unique strengths lie in its self-updating capabilities, its intelligent caching for rapid rebuilds, and its deep integration with the Git-based workflows that are common in computational research. We also note other modern typesetting systems like Typst Typst2024 and frameworks like Bookdown Xie2016_bookdown, which offer alternative paradigms for scientific document creation. While powerful, Rxiv-Maker continues to evolve. Recent updates have addressed the need for broader collaboration by introducing Word export capabilities, allowing authors to share manuscripts with collaborators who rely on word processors. We also plan to integrate Pandoc pandoc2020 more deeply to enable further multi-format outputs, such as HTML and EPUB, while preserving our commitment to typographic quality. We also intend to enhance the framework’s integration with computational environment managers to further improve the portability and reproducibility of our executable manuscripts across different systems biaflows2024; dl4miceverywhere2024, for which containerisation technologies are also a valid alternative Boettiger2015_docker_reproducibility; gomes2018. By applying the principles of literate programming Knuth1984_literate_programming to the creation of academic papers, Rxiv-Maker helps bridge the gap between computational research and traditional scientific publishing. It fosters a workflow where the manuscript is not merely a report of the research, but a reproducible and verifiable embodiment of it. In an era where the responsibility for typesetting and validation increasingly falls to authors, tools that automate quality control and promote rigour are essential. Rxiv-Maker is our contribution to this movement, offering a framework that empowers researchers to produce transparent, reproducible, and professionally polished manuscripts with confidence and efficiency.
{manuscriptinfo}
This work is licensed under CC BY 4.0.
{data}
The arXiv monthly submission data used in this article is available at https://arxiv.org/stats/monthly_submissions. Preprint submissions data across different hosting platforms is available at https://github.com/esperr/pubmed-by-year. The source code and data for the figures in this article are available at https://github.com/HenriquesLab/rxiv-maker.
{code}
The Rxiv-Maker computational framework is available at https://github.com/HenriquesLab/rxiv-maker. The companion Visual Studio Code extension is at https://github.com/HenriquesLab/vscode-rxiv-maker. For users requiring containerised execution, the docker-rxiv-maker repository provides Docker-based deployment at https://github.com/HenriquesLab/docker-rxiv-maker. All repositories are released under an MIT License.
{contributions}
Bruno M. Saraiva, Guillaume Jacquemet, and Ricardo Henriques conceived the project and designed the framework. António D. Brito beta-tested features and troubleshooted user experience. All authors contributed to writing and reviewing the manuscript.
Acknowledgements.
The authors thank Jeffrey Perkel for feedback that helped improve the manuscript. B.S. and R.H. acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101001332) (to R.H.) and funding from the European Union through the Horizon Europe program (AI4LIFE project with grant agreement 101057970-AI4LIFE and RT-SuperES project with grant agreement 101099654-RTSuperES to R.H.). Funded by the European Union. However, the views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them. This work was also supported by a European Molecular Biology Organization (EMBO) installation grant (EMBO-2020-IG-4734 to R.H.), a Chan Zuckerberg Initiative Visual Proteomics Grant (vpi-0000000044 with https://doi.org/10.37921/743590vtudfp to R.H.), and a Chan Zuckerberg Initiative Essential Open Source Software for Science (EOSS6-0000000260). This study was supported by the Academy of Finland (no. 338537 to G.J.), the Sigrid Juselius Foundation (to G.J.), the Cancer Society of Finland (SyƶpƤjƤrjestƶt, to G.J.), and the Solutions for Health strategic funding to ƅbo Akademi University (to G.J.). This research was supported by the InFLAMES Flagship Program of the Academy of Finland (decision no. 337531).
[…]
Supplementary Information
[…]
{stable*}
[ht] Format Input Extension Processing Method Output Formats Quality Use Case Mermaid Diagrams .mmd Mermaid CLI SVG, PNG, PDF Vector/Raster Flowcharts, architectures Python and R Figures .py, .R Script execution PNG, PDF, SVG Publication Data visualisation Static Images .png, .jpg, .svg Direct inclusion Same format Original Photographs, logos LaTeX Graphics .tex, .tikz LaTeX compilation PDF Vector Mathematical diagrams Data Files .csv, .json, .xlsx Python and R processing Via scripts Computed Raw data integration Supported Figure Generation Methods. Overview of the framework’s figure processing capabilities, demonstrating support for both static and dynamic content generation with an emphasis on reproducible computational graphics.
{stable*}
[ht] Tool Type Markdown Primary Use Case Key Strengths Open Source Rxiv-Maker Pipeline Excellent Reproducible preprints Local-first execution, automated caching, rich CLI Yes Overleaf Overleaf2024 Web Editor Limited Collaborative LaTeX Real-time collaboration, rich templates, cloud-based Freemium Quarto Quarto2024 Publisher Native Multi-format publishing Polyglot support, multiple outputs, scientific focus Yes Manubot himmelstein2019 Collaborative Native Version-controlled writing Automated citations, transparent collaboration, Git-based Yes Pandoc pandoc2020 Converter Excellent Format conversion Universal format support, extensible filters Yes Typst Typst2024 Typesetter Good Modern typesetting Fast compilation, modern syntax, growing ecosystem Yes Bookdown Xie2016_bookdown Publisher R Markdown Academic books Cross-references, multiple formats Yes Direct LaTeX Typesetter None Traditional publishing Full control, established workflows, mature ecosystem Yes Comparison of Manuscript Preparation Tools. This comparison positions each tool within the scientific publishing ecosystem. Rxiv-Maker specialises in reproducible preprint workflows with local-first execution and developer-centric features. Other tools address distinct needs, such as real-time collaborative editing (Overleaf), multi-format output (Quarto), or version-controlled writing (Manubot).
[…]
A Getting Started Tutorial
To help new users begin with Rxiv-Maker, this tutorial provides a simple walkthrough for creating a minimal manuscript. The process starts with initialising a new project, which sets up the necessary directory structure and configuration files.
First, open a terminal and run the command rxiv init my_new_paper. This creates a new directory named my_new_paper containing the essential files: 00_CONFIG.yml for manuscript metadata, 01_MAIN.md for the main text, 02_SUPPLEMENTARY_INFO.md for supplementary content, and 03_REFERENCES.bib for the bibliography. The system also creates subdirectories for FIGURES and DATA.
Next, you can edit the 01_MAIN.md file to add your text. You can use standard Markdown for formatting, such as ## for headings and * for italics. To add a citation, simply use the @key syntax, where key corresponds to an entry in your .bib file. For example, to cite the original literate programming paper, you would write [@Knuth1984_literate_programming].
To build the PDF, navigate into the project directory (cd my_new_paper) and run the command rxiv pdf. The framework will process your files, execute any embedded code, and generate a professionally typeset PDF in the output directory. If you encounter any issues, running rxiv validate provides a detailed check of your manuscript’s integrity, flagging common problems like missing figures or broken citations.
This straightforward process allows you to get from a blank slate to a compiled PDF, providing a solid foundation that you can then build upon with more advanced features like programmatic figures and tables.
[…]
Rxiv-Maker Markdown Syntax and Advanced LaTeX Integration
This comprehensive reference demonstrates the automated translation system that enables researchers to write in familiar markdown syntax whilst producing professional LaTeX output. The table below showcases both the standard Markdown-to-LaTeX translations and serves as a perfect example of Rxiv-Maker’s {{tex:...}} blocks, which allow direct LaTeX injection for advanced formatting needs.
For complex table structures that require precise control over formatting, multi-column headers, or mathematical notation, Rxiv-Maker’s {{tex:...}} syntax provides full access to LaTeX’s typesetting capabilities. This table itself was created using {{tex:...}} blocks, demonstrating how raw LaTeX can be seamlessly integrated into Markdown documents while maintaining the tex block protection system that prevents markdown processing of LaTeX-specific syntax.
Sup. Table 1: Rxiv-Maker Markdown to LaTeX Translation Reference. Comprehensive mapping of Markdown syntax to corresponding LaTeX commands, demonstrating the automated translation system that enables researchers to write in familiar markup while producing professional typesetting.
Markdown Input LaTeX Output Description
Basic Text Formatting
$\textbf{bold text}$ Bold formatting
$\textit{italic text}$ Italic formatting
__underlined text__ $\underline${underlined text} Underlined formatting for emphasis
$\textbf{\underline{bold and underlined}}$ Nested formatting: bold containing underline
[…]
$\textit{\underline{italic and underlined}}$ Multiple formatting combinations
$\textsubscript{subscript}$ Subscript formatting, e.g., H2O, CO2
$\textsuperscript{superscript}$ Superscript formatting, e.g., E=mc2, xn
Document Structure
# $\section{Header 1}$ Top-level section
## $\subsection{Header 2}$ Second-level section
### $\subsubsection{Header 3}$ Third-level section
Lists
[…]
Links and URLs
[…]
$\url{https://example.com}$ Bare URL
Citations
$\cite{mycitation}$ Single citation
$\cite{himmelstein2019,Overleaf2024}$ Multiple citations
Cross-References
[…]
$\ref{table:label}$ Table reference
[…]
Tables and Figures
[…]
Document Control
<!– comment –> % comment LaTeX comments
[…]
Programmatic Figure Generation
Rxiv-Maker’s figure generation capabilities are designed to ensure a transparent and reproducible connection between your data and your final visualisations. The system supports two main approaches for creating figures programmatically: script-based generation using Python or R, and diagram rendering from text-based descriptions using Mermaid.
For script-based figures, you place your .py or .R scripts in the FIGURES directory. These scripts often leverage powerful plotting libraries such as Matplotlib Hunter2007_matplotlib or Seaborn Waskom2021_seaborn. During compilation, Rxiv-Maker executes these scripts, and any image files they save (e.g., PNG, PDF, SVG) are automatically detected and can be included in your manuscript. This ensures your visualisations are always synchronised with the underlying data and analysis, as a change in one will trigger the regeneration of the other. This is the method used to produce Fig. Methods and Fig. Methods.
For diagrams, such as flowcharts or system architectures, you can use Mermaid Mermaid2023_documentation. You create a .mmd file containing a text-based description of your diagram. The framework uses the Mermaid command-line tool to render this description into a vector or raster image. This allows your diagrams to be version-controlled just like code, making them easy to modify and track over time.
[…]
Mathematical Formula Support
Rxiv-Maker seamlessly integrates mathematical notation by translating Markdown-style expressions into high-quality LaTeX mathematics. This allows you to write complex mathematical content using simple, familiar syntax.
For inline mathematics, you can use single dollar sign delimiters (\$...\$), allowing formulas like $E=mc^{2}$ to be embedded directly within your text. For larger, display-style equations, you can use double dollar signs (\$\$...\$\$) to centre the expression on its own line. For example:
[…]
The framework’s multi-pass translator is designed to protect these mathematical expressions, ensuring they are not accidentally altered during the conversion from Markdown to LaTeX. This robust system supports a wide range of mathematical and statistical notation, from simple symbols to complex multi-line equations, ensuring your quantitative work is always presented clearly and professionally.
[…]
Intelligent Caching and Validation
To accelerate compilation, Rxiv-Maker uses an intelligent caching system that avoids redundant work. It operates by generating a checksum—a unique digital signature—for each figure’s dependencies, including the source script and any input data files. A figure is only regenerated if this checksum changes, meaning you can recompile your manuscript rapidly during writing, as only the modified components are rebuilt. This saves considerable time without sacrificing the reproducibility of the final document.
Complementing this is a powerful validation framework that acts as a quality-control mechanism. Running rxiv validate performs a multi-level check of your manuscript. Before compilation, it looks for missing figure files, broken cross-references, and malformed bibliography entries. During compilation, it parses LaTeX logs to provide clear, understandable error messages. After compilation, it even performs a lightweight scan of the PDF to flag potential rendering issues. This ensures your manuscript is technically sound at every stage.
{sfigure*}
[p!] [Uncaptioned image] The growth of preprint submissions on the arXiv server (1991-2025). This figure was generated from public arXiv statistics using a Python script executed by the Rxiv-Maker pipeline, demonstrating reproducible, data-driven visualisation.
{sfigure*}
[p!] [Uncaptioned image] Preprint Submission Trends Across Multiple Servers (2018-2025). This figure, showing preprints indexed by PubMed from major repositories, was generated from public data PubMedByYear2025 using a reproducible R script within the Rxiv-Maker pipeline.
{sfigure*}
[p!] [Uncaptioned image] The Rxiv-Maker Visual Studio Code Extension. The extension enhances the authoring experience by providing syntax highlighting for extended Markdown, autocompletion for citation keys (e.g., @Knuth...), and real-time validation to catch errors as you type.